OpenStudy (anonymous):

Please help me with this calculus problem! 1.a) Differentiate each function: i) (-pi/2)sin(20x-pi) =10picos(20x) ii) sin^2(cos^3(x)+tan(x) =2cos(cos^3(x)+tan(x))(sin(cos^3(x)+tan(x))(sec^2(x)+3(-sin(x))cos^2(x)) iii) 4x^2sin^2(6x^2-2) =4(2x)sin^2(2-6x^2)-24x^3sin(2-6x^2)cos(2-6x^2)) 1.b) Determine f''(x) where f(x)=cot(4x) and f(x)=sin(x) times tan(x) I need help with this one.

OpenStudy (anonymous):

It looks like you have the answers. What are you asking?

OpenStudy (anonymous):

For part b)

OpenStudy (anonymous):

I dont know what to do

OpenStudy (anonymous):

You have two different f(x) equations. You said f(x)=cot(4x) and f(x)=sin(x) not sure what you want with this. Do you want the second derivative of both?

OpenStudy (anonymous):

f(x)=cot(4x) f'(x) = - 4 csc² 4x

OpenStudy (anonymous):

f (x) = sinx tanx f'(x) = cosx tanx + sinx sec²x

OpenStudy (anonymous):

Thanks for your help chlorophyll, its just that, I dont understand what the question is asking and what it has to do with part a)

OpenStudy (anonymous):

All derivative f'(x)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So we're trying to find f''(x)?

OpenStudy (anonymous):

f(x) = (-π/2)sin(20x-π) f'(x) = 20 * (-π/2) cos (20x-π) = -10 π cos (20 x - π )

OpenStudy (anonymous):

So are those answers that I had wrong?

OpenStudy (anonymous):

It seems not quite correctly!

OpenStudy (anonymous):

I checked it with wolframalpha, hmm

OpenStudy (anonymous):

sin² ( cos³ x+tanx ) Is it what you post?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Have you checked wolf fram yet?

OpenStudy (anonymous):

Yes, I checked them all with wolframalpha

OpenStudy (anonymous):

It says you're right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Wolf has little bit different way calculation with manual one!

OpenStudy (anonymous):

But is it still correct?

OpenStudy (anonymous):

Not really, because there's couple of time I helped the askers, wolf answer different from their online solution, so they lost their points!

OpenStudy (anonymous):

Ok disregarding that, how do we do part b)?

OpenStudy (anonymous):

For example, f(x) = (-π/2) sin (20x-π) Let u = 20x - π ( sin u)' = u' cos u

OpenStudy (anonymous):

I've already finished b part first o.o

OpenStudy (anonymous):

it asks for f''(x) not f'(x)

OpenStudy (anonymous):

so just take the next one! f'(x) = - 4 csc² 4x - 4 csc² (4x) = - 4 sin ^-2 ( 4x) => f" (x) = 32 sin ^-3 ( 4x) * cos (4x)

OpenStudy (anonymous):

I was just wondering why is it part b) if it has nothing to do with part a)?

OpenStudy (anonymous):

It's all individually questions, there's no related at all.

OpenStudy (anonymous):

ok, thanks for your help :)